RANK TESTS BASED ON EXCEEDING OBSERVATIONS

EUGENIA STOIMENOVA

Institute of Mathematics, Bulgarian Academy of Sciences, Acad. G. Bontchev str. bl.8, 1113 Sofia, Bulgaria

(Received February 23, 1998; revised November 5, 1998)

Abstract. Rank tests based on the maximum number of exceeding observations for several standard nonparametric hypotheses are proposed. An approach to constructing nonparametric rank tests via metrics on the permutation group is used. The test statistics are based on a metric induced by Chebyshev’s norm.

Key words and phrases: Induced rank test statistic, location problem, multi-sample problems, partial rankings, metric on permutations, invariance, Chebyshev’s metric, subgroups of permutations.

————————-